# STRRINGS: STReams in Residual Images of Nearby GalaxieS

### v1

@@ Abstract
Tidal features from galaxy mergers, particularly stellar streams, offer valuable insights into galaxy assembly and dark matter halo properties. This paper aims to identify a large sample of nearby stellar streams suitable for detailed modelling and comparison with simulations to enable population-level constraints on halo properties. We visually inspect and compile a tidal feature catalogue for galaxies with redshift from the Siena Galaxy Atlas 2020 using original, model, and residual images from the DESI Legacy Imaging Surveys. Residual images, produced by subtracting models of all sources, enhance the detectability of faint asymmetries such as tidal features. We find that of galaxies host tidal features, more frequently around early-type than late-type galaxies. The tidal feature fraction increases with stellar mass, from at M⊙ to at M⊙. From this, we present the first release of STRRINGS: STReams in Residual Images of Nearby GalaxieS, a subsample of 35 galaxies with long, narrow streams suitable for modelling. We segment these streams and derive their geometry, surface brightness, colours, and stellar masses. The median -band surface brightness is 26.8 mag arcsec-2, reaching 27.5 mag arcsec-2 for the faintest stream. Mass ratios are consistent with minor mergers, and we identify five potential dwarf galaxy progenitors. Our streams are typically longer (median 124 kpc) than the literature, with comparable widths. Stream mass correlates with length and colour, and wider streams lie at larger galactocentric radii. STRRINGS will be expanded and used to constrain halo properties in future work.
@@
The advent of photometric (e.g., SDSS, 2MASS, DES, York et al., 2000; Skrutskie et al., 2006; Dark Energy Survey Collaboration et al., 2016), spectroscopic (LAMOST, APOGEE, Cui et al., 2012; Majewski et al., 2017) and astrometric surveys (Gaia, Gaia Collaboration et al., 2016) allowed for a precise characterisation of the streams system of the Milky Way (MW) (e.g., see review by Bonaca & Price-Whelan, 2025). When the 6D phase-space information is available, tight constraints can be put on the galaxy mass and dark matter halo shape (e.g., Malhan & Ibata, 2019; Dodd et al., 2022; Koposov et al., 2010, 2023). Constraints can also be derived using reduced-dimensional datasets. For example, Erkal et al. (2019) utilised 5D information on the Orphan stream (on-sky position, distance, and proper motions) and showed that the observed misalignment between stellar velocity vectors and the stream track could be attributed to the significant gravitational influence of the Large Magellanic Cloud (LMC). They used this 5D information and simultaneously fitted the Milky Way and LMC potential to provide a mass estimate of the LMC. Fardal et al. (2013) used radial velocities and distance measurements of the Giant Stellar Stream to infer the mass of M31’s halo.
@@
Over the past decade, the search for stellar streams and tidal features around external galaxies has rapidly advanced. Their detection was long challenged by their intrinsically low surface brightness (LSB), but has become increasingly possible due to progress in observational strategies, CCD advancements, and improved data processing pipelines. Searches for LSB tidal features has been mostly conducted visually (e.g., Schweizer & Seitzer, 1992; Tal et al., 2009; Kaviraj, 2010; Sheen et al., 2012; Atkinson et al., 2013; Casteels et al., 2013; Morales et al., 2018; Hood et al., 2018; Müller et al., 2019; Bílek et al., 2020; Yoon & Lim, 2020; Sola et al., 2022, 2025; Huang & Fan, 2022; Jackson et al., 2023; Martínez-Delgado et al., 2023; Skryabina et al., 2024; Miró-Carretero et al., 2024b; Yoon et al., 2024; Pippert et al., 2025). The advent of surveys covering thousands of square degrees such as Euclid (Euclid Collaboration et al., 2022) or Rubin/LSST (Martin et al., 2022) will push even further the limit of the faint Universe. Automated methods (e.g., Conselice et al., 2003; Abraham et al., 2003; Lotz et al., 2004; Pawlik et al., 2016; Kado-Fong et al., 2018; Mantha et al., 2019; Hendel et al., 2019) will become necessary to deal with the huge amount of LSB-compliant images. Another avenue in the search for tidal features is machine learning, to quickly process millions of images, whether using supervised learning (such as convolutional neural networks) that require training sets of pre-annotated images (e.g., Pearson et al., 2019a; Bickley et al., 2021; Walmsley et al., 2022a; Domínguez Sánchez et al., 2023; Gordon et al., 2024), semi-supervised or unsupervised learning (e.g., Desmons et al., 2024).
@@ 2.1 DESI Legacy Imaging Surveys
In addition to images, DESI-LS provides source catalogues that are generated using an inference-based forward-modelling approach, the Tractor (Lang et al., 2016). The processing pipeline, legacypipe222https://github.com/legacysurvey/legacypipe (Dey et al., 2019), starts by extracting sources in the individual bands and stacked images. Each source is then modelled by the Tractor as either a point source, a round exponential galaxy with a variable radius, a deVaucouleurs profile, an exponential profile or a Sersic profile. The best-fit model (based on a likelihood estimation) is kept, and forced photometry is applied to the mid-infrared bands. This produces catalogues of object position, fluxes and colours measured in a consistent way across all DESI-LS images. Therefore, for each image, a model is created and a residual image is obtained by subtracting the model from the image.
Residual images play a central role in our detection of faint tidal features. Because standard modelling algorithms are optimised for compact, high surface brightness structures, they typically fail to capture the extended, diffuse light from LSB features. As a result, these structures are not included in the model and remain visible in the residuals. In an ideal case, residual images should be flat, with any significant deviation indicating the presence of unmodelled components such as tidal features or other faint structures. An illustration of the legacypipe pipeline applied to an image of a galaxy with a stream is shown in Figure 1. However, any inaccuracies in the modelling of stars, galaxies, or internal galactic structures can introduce artefacts in the residual images. High residuals also commonly appear around bright stars, in the extended stellar haloes of galaxies. To mitigate these issues, tidal feature classification must be based on the simultaneous inspection of the original, model, and residual images in order to distinguish real features from artefacts. Stretching the original image would obviously make tidal feature identification easier, but they appear more clearly in the residuals.
@@ 2.3 Galaxy sample
In this paper, we restrict our study to nearby galaxies with redshifts z0.02, corresponding to a maximum distance of approximately 85 Mpc666Cosmology calculator (Wright, 2006), https://astro.ucla.edu/~wright/CosmoCalc.html, assuming km s-1Mpc-1, , . This arbitrary distance threshold allows us to build a large statistical sample; however, we plan to extend the search for streams and tidal features to more distant galaxies in future work. Our redshift criteria limits the sample to individual galaxies. Some are located in galaxy groups, in which case we consider the group as a single object. The definition of a group is based on a projected angular separation between galaxies smaller than 10′ (Moustakas et al., 2023), without requiring them to be physically bounded. This results in a final sample of objects. The mean angular size (D25_LEDA) of galaxies with is 47′′.
@@ 2.4 Images
For each galaxy, we download the already available PNG montage of the colour image, model and residuals777Available at https://portal.nersc.gov/project/cosmo/data/sga/2020/html/.. The field-of-view corresponds to three times the size of the galaxy888The galaxy size is taken in the legacypipe pipeline as D25, the approximate major-axis diameter at the 25 mag arcsec-2 optical SB isophote.. To enhance the visibility of residuals, we convert the PNG images to black and white, and we clip the values to only show the ones between the 35% and 90% percentiles of the pixel distribution. These images, illustrated in Figure 12, are used to perform the classification through visual inspection. While their availability significantly accelerates the data processing workflow, this comes at the expense of a limited field of view of three times the galaxy size, which may reduce the number of detectable features (see Section 6.2). This trade-off was intentionally made to enable the rapid inspection of thousands of images, as our focus is primarily on tidal features located in the immediate vicinity of the target galaxies.
@@ 3.1 Visual classification
For each galaxy, we start by checking if the galaxy has a feature of any type or not: in case of ties, the galaxy is considered as having a feature. We then remove the ‘No feature’ from the votes and we determine majority feature, i.e. the label that occurs the most frequently. In case of ties between several labels, we always favour streams before any other label, then tidal tails, then shells, then plumes, then disturbed/asymmetric halo, then grand-design spiral, then curiosity. The confidence score of the majority feature is the number of votes for that feature divided by the total number of votes.
@@ 4 Tidal features catalogue
Table 2 summarises the fraction of galaxies with a given type of feature, with their standard errors on proportions, as a function of the morphological type. The majority of the galaxies () do not display any peculiarity, while have streams, have tails, have shells, have plumes. Hence, of the galaxies host any type of tidal feature (stream, tail, plume or shell), and display disturbed or asymmetric stellar haloes. An additional are flagged as being grand-design spirals and appear as curiosities.
When considering ETGs and LTGs separately, of ETGs host tidal features against for LTGs, which is a statistically significant difference at a 5% confidence level according to a Z-test on proportions. Shells are more frequent around ETGs than LTGs, which is statistically confirmed by a Z-test. Tails seem to be more present around LTGs (the conditions of application of the Z-test were not verified so it could not be used). The fraction of galaxies with streams is similar for ETGs and LTGs for a galaxy mass below . However, for higher masses, ETGs have a statistically significantly higher stream fraction than LTGs.
@@
Finally, we investigate the impact of the environment, characterised by the group multiplicity from the SGA-2020 catalogue, on the tidal feature fraction. We separate galaxies between isolated (i.e., a group multiplicity of one) and in groups (i.e., a multiplicity higher or equal to two), but we do not make a difference between galaxy groups and clusters. The majority (92.3%) of the galaxies in our sample are isolated. The tidal feature fraction is much higher () for groups than for isolated galaxies (), which is statistically significant according to a Z-test. In groups, the fraction of streams (tails) is (), compared to () for isolated galaxies. Higher fractions in galaxy groups are expected due to the proximity of companion galaxies, hence potential mergers, compared to isolated galaxies. These results on the tidal feature catalogue will be discussed and compared to literature values in Section 6.1.
@@ 5.1 STRRINGS sample selection
We select the 861 galaxies where the dominant feature is classified as ‘streams’. Four classifiers then re-examined these images using the same classification tool, now with three refined options: (1) ‘Not streams’ (the feature is not a stream or is overshadowed by multiple other features), (2) ‘Stream not selected for modelling’ (due to insufficient curvature for our method, excessive brightness, or central galaxy perturbation), and (3) ‘Stream selected for modelling’. We emphasise that this classification is designed for our specific modelling goals, other studies may adopt different criteria depending on their scientific objectives or modelling frameworks. Our final STRRINGS sample includes only objects where all classifiers agreed on a stream selected for modelling, resulting in 35 galaxies and their streams, which are marked with a (*) in Table 1. We then annotate these streams using Jafar, carefully delineating their contours primarily from the residual image while simultaneously checking with the original image.
@@ 5.2 STRRINGS stream properties
We retrieve quantitative measurements of the segmented streams from the database, and we compute the coordinates, length and width of the track. Using the coordinates of the contours of the annotations, we create masks of the streams and apply them to the residual cutout images from the DESI-LS viewer to derive their surface brightness (SB) and colours. An illustration of the stream segmentation and track is shown in Figure 3, while several of our STRRINGS streams are shown in Figure 4. The images and segmentations for all the STRRINGS sample is shown in Figure 17. We provide in Appendix B the summary of the geometric and photometric properties of the streams in Table 4 and the coordinates of the contours of the annotations in a format readable by SAOImageDS9 and Aladin. We checked that the photometry estimated from the residuals is comparable to the one obtained from the original images in Appendix C.
@@ 5.2.1 Surface brightness and colours
The surface brightness of each stream is not uniform but varies along its length. Additional variability arises from the measurement method, whether calculated per pixel or within small aperture boxes. In this study, we report the median SB values, as they provide the most representative estimate and best reflect the central tendency of the distribution. Figure 5 presents the kernel density estimates111111In this work, we plot kernel density estimates instead of histograms because of their smoother representation of distributions. We use the scipy.stats.gaussian_kde function, which applies Gaussian kernels, with the bandwidth determined using the default ‘scott’ method. of the distributions of the streams median SB and colours. The streams have a median surface brightness of 26.4, 26.8 and 25.1 mag arcsec-2 in the and -band, respectively. The faintest feature reaches median SB of 27.4, 27.5 and 25.4 mag arcsec-2 in , and -band, respectively (where the SB limits are 28.5, 29 and 27.4 mag arcsec-2, respectively).
The distribution of the median and -band SB values are similar to the mean SB values reported by Martínez-Delgado et al. (2023) (24 streams), Miró-Carretero et al. (2023) (22 streams) and Miró-Carretero et al. (2024b) (63 streams) who utilised DESI-LS images reprocessed to avoid oversubtraction of the background, leading to a few fainter streams. A thorough comparison of the streams in common between STRRINGS and these literature works will be presented in Sections 6.2 and B.2. The -band are similar to the ones reported by Skryabina et al. (2024) for their 9 streams (measured on the SDSS Stripe 82 images, with a depth of 28.6 mag arcsec-2)121212http://research.iac.es/proyecto/stripe82/. The fainter 100 streams in Sola et al. (2025) are found in deeper CFHT images (with a depth of about 29 mag arcsec-2) for galaxies with a maximum distance of 45 Mpc. The 0.8 mag discrepancy in the -band between our values and those of Martínez-Delgado et al. (2023) and Miró-Carretero et al. (2024b) could be due to their more refined background re-estimation. This implies that our and colours are systematically offset from theirs and not reliable. The streams median colour of 0.5 is in agreement with the literature.

### v2

@@ Abstract
Tidal features from galaxy mergers, particularly stellar streams, offer valuable insights into galaxy assembly and dark matter halo properties. This paper aims to identify a large sample of nearby stellar streams suitable for detailed modelling and comparison with simulations to enable population-level constraints on halo properties. We visually inspect and compile a tidal feature catalogue for galaxies with redshift from the Siena Galaxy Atlas 2020 using original, model, and residual images from the DESI Legacy Imaging Surveys. Residual images, produced by subtracting models of all sources, enhance the detectability of faint asymmetries such as tidal features. We find that of galaxies host detectable tidal features, more frequently around early-type than late-type galaxies. The tidal feature fraction increases with stellar mass, from at M⊙ to at M⊙. From this, we present the first release of STRRINGS: STReams in Residual Images of Nearby GalaxieS, a subsample of 35 galaxies with long, narrow streams suitable for modelling. We segment these streams and derive their geometry, surface brightness, colours, and stellar masses. The median -band surface brightness is 26.8 mag arcsec-2, reaching 27.5 mag arcsec-2 for the faintest stream. Mass ratios are consistent with minor mergers, and we identify five potential dwarf galaxy progenitors. Our streams are typically longer (median 124 kpc) than the literature, with comparable widths. Stream mass correlates with length and colour, and wider streams lie at larger galactocentric radii. STRRINGS will be expanded and used to constrain halo properties in future work.
@@
The advent of photometric (e.g., SDSS, 2MASS, DES, York et al., 2000; Skrutskie et al., 2006; Dark Energy Survey Collaboration et al., 2016), spectroscopic (LAMOST, APOGEE, Cui et al., 2012; Majewski et al., 2017) and astrometric surveys (Gaia, Gaia Collaboration et al., 2016) allowed for a precise characterisation of the stream system of the Milky Way (MW) (e.g., see review by Bonaca & Price-Whelan, 2025). When the 6D phase-space information is available, tight constraints can be put on the galaxy mass and dark matter halo shape (e.g., Malhan & Ibata, 2019; Dodd et al., 2022; Koposov et al., 2010, 2023). Constraints can also be derived using reduced-dimensional datasets. For example, Erkal et al. (2019) utilised 5D information on the Orphan stream (on-sky position, distance, and proper motions) and showed that the observed misalignment between stellar velocity vectors and the stream track could be attributed to the significant gravitational influence of the Large Magellanic Cloud (LMC). They used this 5D information and simultaneously fitted the Milky Way and LMC potential to provide a mass estimate of the LMC. Fardal et al. (2013) used radial velocities and distance measurements of the Giant Stellar Stream to infer the mass of M31’s halo.
@@
Over the past decade, the search for stellar streams and tidal features around external galaxies has rapidly advanced. Their detection was long challenged by their intrinsically low surface brightness (LSB), but has become increasingly possible due to progress in observational strategies, CCD advancements, and improved data processing pipelines. Searches for LSB tidal features has been mostly conducted visually (e.g., Schweizer & Seitzer, 1992; Tal et al., 2009; Kaviraj, 2010; Sheen et al., 2012; Atkinson et al., 2013; Casteels et al., 2013; Morales et al., 2018; Hood et al., 2018; Müller et al., 2019; Bílek et al., 2020; Yoon & Lim, 2020; Sola et al., 2022, 2025; Huang & Fan, 2022; Jackson et al., 2023; Martínez-Delgado et al., 2023; Skryabina et al., 2024; Miró-Carretero et al., 2024b; Yoon et al., 2024; Pippert et al., 2025). The advent of surveys covering thousands of square degrees such as Euclid (Euclid Collaboration et al., 2022) or Rubin/LSST (Martin et al., 2022) will push even further the limit of the faint Universe. Automated methods (e.g., Conselice et al., 2003; Abraham et al., 2003; Lotz et al., 2004; Pawlik et al., 2016; Kado-Fong et al., 2018; Mantha et al., 2019; Hendel et al., 2019) will become necessary to deal with the huge amount of LSB-compliant images. Another avenue in the search for tidal features is machine learning, to quickly process millions of images, whether using supervised learning (such as convolutional neural networks) that require training sets of pre-annotated images (e.g., Walmsley et al., 2019; Pearson et al., 2019a; Bickley et al., 2021; Walmsley et al., 2022a; Domínguez Sánchez et al., 2023; Gordon et al., 2024), semi-supervised or unsupervised learning (e.g., Desmons et al., 2024).
@@ 2.1 DESI Legacy Imaging Surveys
In addition to images, DESI-LS provides source catalogues that are generated using an inference-based forward-modelling approach, the Tractor (Lang et al., 2016). The processing pipeline, legacypipe222https://github.com/legacysurvey/legacypipe (Dey et al., 2019), starts by extracting sources in the individual bands and stacked images. Each source is then modelled by the Tractor as either a point source, a de Vaucouleurs profile, an exponential profile, a Sérsic profile or a round exponential galaxy model (i.e. with zero ellipticity) with a variable half-light radius. The best-fit model (based on a likelihood estimation) is kept, and forced photometry is applied to the mid-infrared bands. This produces catalogues of object position, fluxes and colours measured in a consistent way across all DESI-LS images. Therefore, for each image, a model is created and a residual image is obtained by subtracting the model from the image.
Residual images play a central role in our detection of faint tidal features. Because standard modelling algorithms are optimised for compact, high surface brightness structures, they typically fail to capture the extended, diffuse light from LSB features. As a result, these structures are not included in the model and remain visible in the residuals. In an ideal case, residual images should be flat, with any significant deviation indicating the presence of unmodelled components such as tidal features or other faint structures. An illustration of the legacypipe pipeline applied to an image of a galaxy with a stream is shown in Figure 1. For better visualisation, a black-and-white residual image can be produced by computing the median of the residual images in g, r, and z, then clipping the pixel distribution. However, any inaccuracies in the modelling of stars, galaxies, or internal galactic structures can introduce artefacts in the residual images. High residuals also commonly appear around bright stars, in the extended stellar haloes of galaxies. To mitigate these issues, tidal feature classification must be based on the simultaneous inspection of the original, model, and residual images in order to distinguish real features from artefacts. Stretching the original image would obviously make tidal feature identification easier, but they appear more clearly in the residuals.
@@ 2.3 Galaxy sample
In this paper, we restrict our study to nearby galaxies with redshifts z0.02, corresponding to a maximum distance of approximately 85 Mpc666Cosmology calculator (Wright, 2006), https://astro.ucla.edu/~wright/CosmoCalc.html, assuming km s-1Mpc-1, , . This arbitrary distance threshold allows us to build a large statistical sample; however, we plan to extend the search for streams and tidal features to more distant galaxies in future work. Our redshift criteria limits the sample to individual galaxies. Some are located in galaxy groups, in which case we only consider the central, primary galaxy and remove from our sample the other galaxies. The definition of a group is based on a projected angular separation between galaxies smaller than 10′ (Moustakas et al., 2023), without requiring them to be physically bounded. This results in a final sample of objects. The mean angular size (D25_LEDA) of galaxies with is 47′′.
@@ 2.4 Images
For each galaxy, we download the already available montage PNG images 777Available at https://portal.nersc.gov/project/cosmo/data/sga/2020/html/. which are 3-channel colour composites of the original, model, and residual data produced by the legacypipe pipeline. Each channel is stored as integer values on a 0–255 scale. The field-of-view corresponds to three times the size of the galaxy888The galaxy size is taken in the legacypipe pipeline as D25, the approximate major-axis diameter at the 25 mag arcsec-2 optical SB isophote.. From these colour images, we compute the median of the residuals in g, r, and z, resulting in a grayscale image. To enhance faint features, we calculate the 35th–90th percentile range of the pixel distribution in this median residual image and clip the pixel values accordingly.
These images, illustrated in Figure 13, are used to perform the classification through visual inspection. While their availability significantly accelerates the data processing workflow, this comes at the expense of a limited field of view of three times the galaxy size, which may reduce the number of detectable features (see Section 6.2). This trade-off was intentionally made to enable the rapid inspection of thousands of images, as our focus is primarily on tidal features located in the immediate vicinity of the target galaxies.
@@ 3.1 Visual classification
For each galaxy, we start by checking if the galaxy has a feature of any type or not: in case of ties, the galaxy is considered as having a feature. Each galaxy is assigned a single label, as the categories are mutually exclusive. We then remove the ‘No feature’ from the votes and we determine majority feature, i.e. the label that occurs the most frequently. In case of ties between several labels, we always favour streams before any other label, then tidal tails, then shells, then plumes, then disturbed/asymmetric halo, then grand-design spiral, then curiosity. This ordering reflects both diagnostic power and clarity of identification. Streams are given the highest priority, with tidal tails ranked second due to the potential ambiguity in distinguishing them from streams. Shells are placed third, as their presence provides a strong indication of tidal interactions, most likely associated with radial mergers. Plumes are ranked next, as they are less well-defined than tails. Disturbed stellar haloes are prioritised after plumes, since they indicate recent interactions without any clearly recognisable substructure. The two optional categories, grand-design spirals and curiosity, are placed last. Finally, we define the confidence score of the majority feature as the number of votes for that feature divided by the total number of votes.
@@ 4 Tidal features catalogue
Table 2 summarises the fraction of galaxies with a given type of feature, with their standard errors on proportions, as a function of the morphological type. The majority of the galaxies () do not display any peculiarity, while have streams, have tails, have shells, have plumes. Hence, of the galaxies host any type of detectable tidal feature (stream, tail, plume or shell), and display disturbed or asymmetric stellar haloes. An additional are flagged as being grand-design spirals and appear as curiosities.
When considering ETGs and LTGs separately, of ETGs host tidal features against for LTGs, which is a statistically significant difference at a 5% confidence level according to a Z-test on proportions. Shells are more frequent around ETGs than LTGs, which is statistically confirmed by a Z-test. Tails seem to be more present around LTGs (the conditions of application of the Z-test were not verified so it could not be used111111A Z-test for comparing two proportions can be applied when the samples are independent, the outcome is binomial, the data are randomly sampled, and the sample sizes are large enough for the normal approximation to be valid (typically np5 and n(1-p)5 for each group, where n is the sample size and p the proportion). This last condition was not met here.). The fraction of galaxies with streams is similar for ETGs and LTGs for a galaxy mass below . However, for higher masses, ETGs have a statistically significantly higher stream fraction than LTGs.
@@
Finally, we examine the role of environment, quantified by group multiplicity in the SGA-2020 catalogue, on the tidal feature fraction. We emphasise that our analysis considers only the central, primary galaxy of each group and does not include tidal features around satellite members. Galaxies are divided into isolated systems (group multiplicity of one) and central galaxies in groups (multiplicity 2), without distinguishing between groups and clusters. Most galaxies in our sample (92.3%) are isolated. The overall tidal feature fraction is significantly higher for central galaxies in groups () than for isolated galaxies (), a difference confirmed by a Z-test. In groups, streams and tails are found in and of galaxies, respectively, compared to only and for isolated systems. This enhancement cannot be attributed solely to stellar mass, as at fixed host mass the tidal feature fraction for group centrals remains two to three times higher than for isolated galaxies. Such elevated fractions in dense environments are expected, given the greater availability of nearby companions and the higher likelihood of mergers. These results on the tidal feature catalogue will be discussed and compared to literature values in Section 6.1.
@@ 5.1 STRRINGS sample selection
We select the 861 galaxies where the dominant feature is classified as ‘streams’. Four classifiers then re-examined these images using the same classification tool, now with three refined options: (1) ‘Not streams’ (the feature is not a stream or is overshadowed by multiple other features), (2) ‘Stream not selected for modelling’ (due to insufficient curvature for our method, excessive brightness, or central galaxy perturbation), and (3) ‘Stream selected for modelling’. We emphasise that this classification is designed for our specific modelling goals, other studies may adopt different criteria depending on their scientific objectives or modelling frameworks.
These selection criteria inevitably introduce biases in the statistical properties of the galaxies found to host tidal streams. By excluding systems with clearly disturbed morphology, we preferentially select dynamically quiet galaxies and avoid recent or ongoing major mergers. Our focus on thin, curved, and faint streams further biases the sample toward minor mergers with low mass ratios, as these produce streams that are best suited for our modelling assumptions. In addition, stream detectability is enhanced in galaxies viewed closer to edge-on, where the disc can be more easily separated from surrounding substructure, which may favour higher inclination hosts. As a result, the STRRINGS sample is not representative of the full population of galaxies with streams, but rather of those hosting streams that can be robustly modelled within our framework.
Our final STRRINGS sample includes only objects where all classifiers agreed on a stream selected for modelling, resulting in 35 galaxies and their streams, which are marked with a (*) in Table 1. We then annotate these streams using Jafar, carefully delineating their contours primarily from the residual image while simultaneously checking with the original image.
@@ 5.2 STRRINGS stream properties
We retrieve quantitative measurements of the segmented streams from the database, and we compute the coordinates, length and width of the track. Using the coordinates of the contours of the annotations, we create masks of the streams and apply them to the residual cutout images from the DESI-LS viewer to derive their surface brightness (SB) and colours. An illustration of the stream segmentation and track is presented in Figure 3, while Figure 4 shows the Cartesian projection of all the streams, centred around their respective host galaxy. Several of our STRRINGS streams in residual images are illustrated in Figure 5. The images and segmentations for all the STRRINGS sample is shown in Figure 18. We provide in Appendix B the summary of the geometric and photometric properties of the streams in Table 4 and the coordinates of the contours of the annotations in a format readable by SAOImageDS9 and Aladin. We checked that the photometry estimated from the residuals is comparable to the one obtained from the original images in Appendix C.
@@ 5.2.1 Surface brightness and colours
The surface brightness of each stream is not uniform but varies along its length. Additional variability arises from the measurement method, whether calculated per pixel or within small aperture boxes. In this study, we report the median SB values, as they provide the most representative estimate and best reflect the central tendency of the distribution. Figure 6 presents the kernel density estimates121212In this work, we plot kernel density estimates instead of histograms because of their smoother representation of distributions. We use the scipy.stats.gaussian_kde function, which applies Gaussian kernels, with the bandwidth determined using the default ‘scott’ method. of the distributions of the streams median SB and colours. The streams have a median surface brightness of 26.4, 26.8 and 25.1 mag arcsec-2 in the and -band, respectively. The faintest feature reaches median SB of 27.4, 27.5 and 25.4 mag arcsec-2 in , and -band, respectively (where the SB limits are 28.5, 29 and 27.4 mag arcsec-2, respectively).
The distribution of the median and -band SB values are similar to the mean SB values reported by Martínez-Delgado et al. (2023) (24 streams), Miró-Carretero et al. (2023) (22 streams) and Miró-Carretero et al. (2024b) (63 streams) who utilised DESI-LS images reprocessed to avoid oversubtraction of the background, leading to a few fainter streams. A thorough comparison of the streams in common between STRRINGS and these literature works will be presented in Sections 6.2 and B.2. The -band are similar to the ones reported by Skryabina et al. (2024) for their 9 streams (measured on the SDSS Stripe 82 images, with a depth of 28.6 mag arcsec-2)131313http://research.iac.es/proyecto/stripe82/. The fainter 100 streams in Sola et al. (2025) are found in deeper CFHT images (with a depth of about 29 mag arcsec-2) for galaxies with a maximum distance of 45 Mpc. The 0.8 mag discrepancy in the -band between our values and those of Martínez-Delgado et al. (2023) and Miró-Carretero et al. (2024b) could be due to their more refined background re-estimation. This implies that our and colours are systematically offset from theirs and not reliable. The streams median colour of 0.5 is in agreement with the literature.